OpenStudy (anonymous):

Find the power series for f(x)= 1/(5-x). I have (1/5) SUM (x/5)^ n. Can't I just leave it in this form? Do i have to make it SUM [(x^n/5^n+1)]??

OpenStudy (amistre64):

do long division on it; it creates a power series for it

OpenStudy (amistre64):

1/5+x/5^2+x^2/5^3 .... --------------------- 5-x ) 1 (1-x/5) -------- x/5 (x/5 - x^2/5^2) ------------- x^2/5^2 \[\sum_{n=0}^{inf}\frac{x^n}{5^{n+1}}\]

OpenStudy (amistre64):

you can factor out the 1/5 yes

OpenStudy (amistre64):

im not sure what a simplified form of summations would be tho :) quite frankly, i find simplifications to be duplicitous at best

OpenStudy (anonymous):

It makes the sum look nicer........and thats all. Im not that good at this, so I want to make things as simple as possible.

OpenStudy (amistre64):

remember what a summation is : sum can = ca1 + ca2 + ca3 + ca4 ... we can factor out the c (constant) c sum an = c (a1 + a2 + a3 + a4 ...)

OpenStudy (anonymous):

True. Im leaving it as is then!! Muah ha ha. Thanks....again!

OpenStudy (amistre64):

yw :)